Let α : Z₉* Z₂₇→ Z₂₇ be given by α ((a, b)) = 3b for a ϵ Z₉, b ϵ Z₂₇ and
(a, b) + (c, d) = (a + c, b + d) is given by addition modulo n for each group Zn.
i. Show that α is a homomorphism.
ii. Find Ker(α).
1
Expert's answer
2020-06-29T18:16:30-0400

Given "\alpha : Z₉ \times Z₂₇→ Z₂₇" be given by "\alpha (a, b) = 3b" for "a \u03f5 Z_9, b \u03f5 Z_{27}" and

"(a, b) + (c, d) = (a + c, b + d)" .

(i) Homomorphism: "\alpha((a, b) + (c, d)) = \alpha(a + c, b + d) = 2(b+d) = 2b + 2d"

Also, "\alpha(a,b)+\alpha(c,d) = 2b+2d" .

Hence, "\alpha((a, b) + (c, d)) = \alpha(a, b) + \alpha(c, d)" .

Hence, "\alpha" is a Homomorphism map.

(ii) "ker(\alpha) = \{(a,b): \alpha(a,b)=0 \}"

"\alpha(a,b)=0 \implies 2b = 0 \implies b = 27 n" where "n\in \Z" .

So, "ker(\alpha) = \{(a,27n): n \in \Z\}"